OpenStudy (anonymous):

HELP MEDAL! The first four terms of a sequence are shown below: 6, 2, -2, -6 Which of the following functions best defines this sequence? f(1) = 6, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 2; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) + 2; for n ≥ 1 @satellite73 @mathstudent55 @TuringTest @Nnesha @confluxepic

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

@OOPS don't understand overall question

OpenStudy (anonymous):

6 is the first term , it is f(1) 2 is the second term, it is f(2) -2 is the third term, it is f(3) -6 is the fourth term, it is f(4)

OpenStudy (anonymous):

Now, I pick f(2) Do you agree with me that f(2) = f(1+1)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if I say n =1, so n+1 =2 , right?

OpenStudy (anonymous):

Not sure about that

OpenStudy (anonymous):

now, \(f(\color{red}{n}+1)= f(\color{red}{n})+4\) so, if \(\color{red}{n=1}\), then \(f(\color{red}{1}+1) = f(2) = f(\color{red}{1})+4= 6+4=10\)

OpenStudy (anonymous):

that is the first option, but our f(2) \(\neq 10\), hence, the first option is not the correct one.

OpenStudy (anonymous):

You read and try to understand what I am saying, then test for option 2, 3, 4 to know which one is the correct answer.

OpenStudy (anonymous):

Unless you don't want to study and I waste my time to explain, I don't think it is that hard. :)

OpenStudy (anonymous):

@OOPS I can see it's not hard but i'm having trouble understanding